OpenStudy (maeganr):

solve by graphing x + y = 3

OpenStudy (mathhelp346):

Turn it into slop-intercept form, which is y=mx+b. m is the slope b is the y-intercept

OpenStudy (mathhelp346):

*slope

OpenStudy (maeganr):

So Y= 3x + b?

OpenStudy (mathhelp346):

No

OpenStudy (anonymous):

x + y = 3 we know that the general form of slope is y=mx +c where m is the slope and c is the intercept and so , i am trying to make the given equation to look like the general form => y=(-1)x+3 find the points: